A sample of 36 ATM transactions shows a mean transaction time of 65 seconds with a sample standard deviation of 12 seconds. When testing whether the mean transaction time is greater than 60 seconds, what is the test statistic?

Select one:

0.42

1.50

2.50

3.12

none of the above

Given your answer to question 14, which of the following do we know for sure about the conclusion to an hypothesis test at a 0.005 significance level? Select one:

a correct decision has been made

a Type II error has not been made

a Type I error has not been made

either a Type I error or Type II error has been made
Answer #1

Answer:

Given,

n = 36

mu = 65

s = 12

Null hypothesis Ho : mu <= 60

Alternative hypothesis Ha : mu > 60

consider,

test statistic t = (xbar - mu) / s/sqrt(n)

substitute values

= (65 - 60) / 12/sqrt(36)

= 2.5

degree of freedom = n - 1

= 36 - 1

= 35

Now the critical value of t for right tailed at alpha 0.005 & df 35 t 0.005,35 is 2.7238

Here test statistic > critical value so we reject Ho.

We conclude that there is sufficient evidence to support the claim that the mean exceeds 60 sec